In a Math test, the mean score is 45 and the standard deviation is 4.


Assuming normality, what is the probability that a score picked at random will lie


c. between 35 and 53?

1
Expert's answer
2022-02-23T09:20:50-0500

"\mu=45\\\sigma=4"

Let the random variable "X" represent the math score.

"p(35 \lt x\lt53)=p({35-45\over4}\lt Z\lt {53-45\over4})=p(-2.5\lt Z\lt 2)=\phi(2)-\phi(-2.5)=0.9772-0.0062=0.991"

The probability that a score picked at random will lie between 35 and 53 is 0.991